OpenStudy (dr.gray):

Which number sentence is NOT true? A.|-3| > |1/3| B.|-3|=|-1/3| C.|1/3| < |-3| D.|-1/3| < |-3|

OpenStudy (mathstudent55):

Do you know what absolute value is?

OpenStudy (ninja4535):

The absolute value of a negative number is a positive number. For example |-6|. |-6| equals to 6. With that being so... |-3| > |1/3| is true because |-3| = 3 and it is greater than 1/3 |1/3|<|-3| is true because 1/3 is less than |-3| which equals to 3 |-1/3| < |-3| is true because |-1/3| = 1/3 is less than |-3| which equals to 3 That only leaves |-3| = |-1/3| is not true.